If Penny bought a stock for $80 dollars and could sell it 15 years later for 4 times what she originally paid, what is Penny's return on owning this stock
Answer #1

Total return is given as equal to=4-1=300%

Annual return is given as equal to=4^(1/15)-1=9.6825%